# ShadowNav: Autonomous Global Localization for Lunar Navigation in Darkness

### v2

@@ 3.5.2 Filtering System
Algorithm 2 provides the full outline for the ShadowNav particle filtering algorithm. The algorithm takes as input a Gaussian belief distribution $(\mu_{0},\Sigma_{0})$ assumed for the initial robot position, the number of particles $N_{s}$ to use in the particle filter, and a threshold for the effective number of beliefs $N_{\textrm{eff,thresh}}$ used to trigger resampling (Line 2). The filter is initialized with $N_{s}$ particles drawn from the initial belief distribution and each assigned equal weight (Lines 1-2). Given a new set of crater observations (Line 5), a set of Q-Score measurements is initialized for computing for each individual particle (Line 6). After applying the motion model update to each particle (Line 8), the Q-Score for each updated particle is computed using the procedure from Alg. 1 by comparing against the current measurements (Line 9). As is standard in many particle filtering implementations [Arulampalam et al., 2002], we note that the particle weights are updated in the in $\log$-domain (Line 13) with a normalization step to ensure non-negative weights (Line 11). Next, the number of effective samples $N_{\textrm{eff}}$ at the current iteration is calculated (Line 15). If $N_{\textrm{eff}}$ is below the threshold $N_{\textrm{eff,thresh}}$, then this is seen as an indication of particle filter “degeneracy”, wherein the weights $\{w_{i}^{t}\}$ collapse around a handful of particles. In such a case, a new set of particles are resampled using the systematic resampling scheme (Line 17).
@@
3.6 Simulating Relative Localization
In order to simulate the impact of drift from relative localization, a bias factor is added. This is in addition to the Gaussian noise factor used in our previous work. This bias is a Gaussian distributed magnitude with the mean as the specified percentage drive of relative localization and a constant unit vector direction. The unit vector direction is randomly chose from a uniform distribution of all possible directions at the start of a replay of a trajectory. For example if the relative localization error rate used is 2% then at each step, the bias factor is a Gaussian centered around 2% of the distance traversed in that step. Over the course of many Monte Carlo runs, the bias will be added in many different directions to test for drift in all directions.
@@ 4 Data Collection
In this section, we present details on the datasets generated to validate the efficacy of the proposed ShadowNav approach. We first review the simulation environment used to generated photorealistic stereo data and then review the configuration used to collect nighttime field testing data from the Cinder Lakes, Arizona.
@@ 4.1 Simulation
In this work, we develop a simulation environment using the Blender software, which has become a popular software framework for generating photorealistic images in a Lunar environment[Crues et al., 2023, Cauligi et al., 2023]. Within this simulation, the Hapke lighting model [Hapke, 2002, Hapke, 2012, Schmidt and Bourguignon, 2019] was implemented which approximates the Lunar surface reflectance. This model will simulate the “opposition effect” which leads to a focused point of extreme saturation at a location within an image where the camera ray and light source are at zero phase angle. We implemented the Hapke lighting model using “old highland” parameters of the Moon provided in [Xu et al., 2020]. For our use case, the coherent backscattering opposition effect (CBOE) was left out of our implementation and only the shadow hiding opposition effect (SHOE) was implemented as it dominates most or all lighting calculations.
To obtain a realistic 3D model of the surface geometry, DEMs produced from LROC were utilized.  LROC resolution is typically between $2\text{\,}\mathrm{m}$-$5\text{\,}\mathrm{m}$ which allows for resolving craters of approximately $10\text{\,}\mathrm{m}$. However, as the LROC resolution is insufficient for generating smooth surface imagery, the DEMs from LROC were scaled down to be $0.25\text{\,}\mathrm{m}$ resolution. Crater measurements for simulated data in future discussions were based on this scaled resolution. This scaled DEM was imported into Blender and a surface texture comprised of two scales of fractal Brownian motion. This texture is a natural noise and simulates surface texture features for stereo algorithms to utilize.
@@ 4.1.1 Simulated Dataset
For simulation, we developed two datasets - one with the aim of testing the efficacy of our crater matching algorithm and a second dataset for testing the overall global localization pipeline. The crater matching dataset was created by selecting four craters from the orbital data of diameters between $5\text{\,}\mathrm{m}$ and $20\text{\,}\mathrm{m}$. The simulator was run to generate stereo pairs every $1\text{\,}\mathrm{m}$ at four angles around the crater (0, 90, 180, and 270) and at ranges between $5\text{\,}\mathrm{m}$-$25\text{\,}\mathrm{m}$. Additional datasets were created to test sensitivity to the parameters of the hardware configuration:
@@
  • Light offsets of $0.2\text{\,}\mathrm{m}$ below the camera, inline with the camera and $0.2\text{\,}\mathrm{m}$ above the camera.
  • Light power ranging from low, medium and high.  [^1] As Blender does not accurately model exposure times, the set of “light power” Blender parameters is used to change the brightness of rendered images.
@@ 4.2 Cinder Lakes Apollo Training Area
For the data collection, we performed two types of collections. The first type was geared toward perception tests. This involved moving the data collection rig from approximately $20\text{\,}\mathrm{m}$ away to $5\text{\,}\mathrm{m}$ from the leading crater edge with stops for long exposures roughly every $1\text{\,}\mathrm{m}$. The second type was geared toward localization tests. This entailed moving the data collection rig through the crater field with stops for long exposure approximately every $5\text{\,}\mathrm{m}$-$10\text{\,}\mathrm{m}$. During these stops the rig was directed toward large craters if any were present. In total, in this work, 12 craters ranging from $5\text{\,}\mathrm{m}$-$25\text{\,}\mathrm{m}$ in diameter and three trajectories were collected and used. The three different trajectories are displayed in Figure 5 which shows the GPS path within Google Earth satellite imagery. Of these, two trajectories were captured in the south site and one in the north site. In order to evaluate longer distances between craters, some experiments added drift error as if more distance had been traveled by artificially adding trajectory locations outside of the crater field at the start of a trajectory.
@@ 5.1 Stereo
In order to be able to effectively detect and match crater rims, accurate stereo is required. For detection, outliers in range values or stereo holes could lead to crater rim false positives. Furthermore, inaccurate stereo ranges could lead to poor projections into the orbital map which could cause good image-space detections to match poorly to orbital landmark craters. In order to evaluate stereo performance, we created a simulated dataset of craters with ground truth ranges which was detailed in Section 4.1.
To evaluate stereo performance, three metrics based on [Scharstein et al., 2001] binned across ground truth range values are used. These metrics are:
@@
We note that a crucial difference between the metrics used in [Scharstein et al., 2001] is that range is used rather than disparity as range is what ultimately has the impact on crater matching and we have perfect camera calibrations within simulation.
To validate the impact of hardware system parameters such as light location, camera height, and light power, these parameters were swept and stereo performance evaluated. Light power is used as a proxy for exposure time and light brightness as these are not modelled well in Blender. High light power corresponds to the near range of the image being fully exposed (pixel value of 255). In order to provide additional realism for nighttime imagery, a zero-mean 1.2 pixel intensity sigma Gaussian noise was added to all images and the results of this analysis are in Table 1. In these results, we observe that the optimal configuration for stereo is with a $2.5\text{\,}\mathrm{m}$ camera off ground, light offset $0.2\text{\,}\mathrm{m}$ below the camera, and high light power. Medium light power at a camera height of $2.5\text{\,}\mathrm{m}$, wherein the near range is bright but few of the pixes are fully exposed, demonstrates improved performance but at the slight cost of degraded perofrmance at range. Moreover, although this configuration has robustness to camera height ranges, we observe that a camera height of $1.5\text{\,}\mathrm{m}$ performs slightly worse than other heights and this is attributed to cameras closer to the ground losing vertical resolution with range. Finally, a key takeaway is that placing the camera inline with the camera significantly degrades performance and placing the light above the camera renders performance extremely poor. We found that placing the light below the camera minimizes some of the opposition effect’s impacts and provides some shadowing on the surface for more visible features.
@@ 5.2 Crater Detection and Matching
In order to evaluate crater detection and matching, we utilized real data collected at Cinder Lakes as detailed in Section 4.2. Additional simulated data was collected to evaluate stereo and positional errors as detailed in Secion 4.1. In order to evaluate detection performance we utilize a metric that is percent of the ground truth front rim detected:
@@
,where $px_{\textrm{gt}_{\textrm{matched}}}$ is the number of ground truth crater points in world frame that have a detection that is closest to that ground truth point and $px_{\textrm{gt}}$ is the total number of all ground truth points. This detection does not consider false positives. In order to evaluate crater matching performance, we utilize the Q-Score metric detailed in Section 3.5.1 to evaluate how closely a perception sample aligns to orbital ground truth considering both matches and false positives.
SGBM stereo and discontinuity detection were run across the Cinder Lakes crater dataset to evaluate crater detection and matching performance. Figure 6 shows the results with respect to percent of front rim detected and Q-scores versus range and crater diameter are shown. From these results, we observed that nearly all craters have some level of detection between $5\text{\,}\mathrm{m}$-$15\text{\,}\mathrm{m}$. However, at $15\text{\,}\mathrm{m}$ the $7\text{\,}\mathrm{m}$-crater is no longer detected. Furthermore, the $25\text{\,}\mathrm{m}$ has worse results across most of the set of camera distances from the crater rim. The reduction in performance for smaller and larger craters is expected with our current approach. Since a discontinuity within stereo must be observed and not just an end to stereo detections, a larger crater diameter requires stereo to obtain matches at farther ranges to observe features on the far side. Additionally smaller diameter craters discontinuity gaps are more likely to decrease below threshold levels at increased range. Overall, we observe robust filter performance, with most samples containing a Q-Score of 0.4 or better which means that the average pixel detection is within $2\text{\,}\mathrm{m}$ of the crater rim and indicates low false positives and high accuracy.
@@
In order to further evaluate crater detection, simulated data is evaluated first by comparing crater rim detection and matching with and without Gaussian noise added. Figure 7 contains the results of comparing noise free versus images with zero mean, 1.2 sigma pixel intensity Gaussian noise added. From these results, it is observed that the addition of noise has minimal impact on crater detection performance. Compared with the Cinder Lakes results, however, detection performance is a bit worse at both close to $5\text{\,}\mathrm{m}$ to the crater and beyond $15\text{\,}\mathrm{m}$ from crater rim. The largest difference is in the Q-Score computation with similar crater rim detection performance to Cinder Lakes data. Potential sources of this performance difference include the simulation containing images with a wider field-of-view, a $2.5\text{\,}\mathrm{m}$ camera height compared to $1.5\text{\,}\mathrm{m}$ camera height for Cinder Lakes data, and the inherent simulation versus real differences in lighting, camera exposure handling, and surface textures.
The higher camera height would potentially allow the images to capture the near side of the crater and therefore limit the discontinuity. The larger field-of-view also would allow for more potential false positives to be captured as evidenced by strong crater rim percent detection but degraded Q-Scores at $15\text{\,}\mathrm{m}$-$20\text{\,}\mathrm{m}$ ranges to crater rims.
@@
Figure 7: Crater detection and matching results with an without Gaussian noise added to simulated images on data with a camera height of 2.5m a light offset of 20cm below camera and high exposure. For images with noise added, 0 mean and 1.2 pixel intensity sigma were used. (a) Q score no noise (b) Percent detected no noise (c) Q Score 1.2 sigma noise (d) Percent detected 1.2 sigma noise.
One of the limitations of the above analysis is that all of the craters were mostly centered within the image frame. In this work, we assumed that highly accurate heading estimates are provided from, e.g., a star tracker and allow for centered detections with perfect absolute position. However, upon approaching a crater, there is an expected error in absolute position due to the drift of relative localization. In order to evaluate how these positional errors affect detection performance, data was rendered with a positional offset from the position that the ideal heading was determined from. This data is described in more detail in Section 4.1. Using this data, Figure 9 shows the results of Q-scores and detected rim percent results with a $5\text{\,}\mathrm{m}$ and $10\text{\,}\mathrm{m}$ horizontal shift. Only horizontal errors are plotted because forward/backward errors are captured as a shift along the x-axis in Figure 7. From Figure 9, we observed that $5\text{\,}\mathrm{m}$ horizontal offset has some degredation in crater matching performance but detection stays fairly similar. However, crater matching and detection degrades at less than $10\text{\,}\mathrm{m}$ range due to the crater not being captured within the field-of-view of the camera. Crater matching still degrades around $15\text{\,}\mathrm{m}$, but the degredation is similar to that of no positional offset. Overall crater detection and matching demonstrates strong performance within the $5\text{\,}\mathrm{m}$-$15\text{\,}\mathrm{m}$ range with potential for performance out to $20\text{\,}\mathrm{m}$ detection range.
Figure 9: Crater detection and matching results with positional error on simulated data with a camera height of 2.5m a light offset of 20cm below camera, high brightness, and 1.2 pixel intensity sigma Gaussian noise to image (a) Q score 5m error (b) Percent detected 5m error (c) Q score 10m error (d) Percent detected 10m error.
5.3 Absolute Localization
Instantaneous perception measurements demonstrated strong performance in the previous section, but also presented some issues. Through the use of a particle filter approach outlined in Section 3.4, multiple perception measurements can be fused overtime to correct absolute position after it has drifted. In order to evaluate the performance of the absolute localization both simulation and Cinder Lakes data is used.
5.3.1 Simulation
@@
First, evaluation is performed within simulation to determine the best trajectory to observe a landmark crater. This was accomplished by designing three types of trajectories around the same landmarks. The first trajectory type was a straight trajectory which stops every $10\text{\,}\mathrm{m}$ and travels directly past the crater with the nearest approach approximately $10\text{\,}\mathrm{m}$ from the rim. The second trajectory type was a half survey, which was designed so that observation positions occur 180 degrees around the crater with each observation point approximately $7\text{\,}\mathrm{m}$-$13\text{\,}\mathrm{m}$ from the crater rim. The third trajectory type was similar to half survey but had observation points a full 360 degrees around the crater. Table 2 contains the results for all trajectory types with a 2% error growth and Figure 10 is a histogram of final absolute errors of three different trajectories. For each trajectory, we performed Monte Carlo runs of 30 different tests and each test having a 2% relative localization error. From these results, we observed that the half survey and the full survey approaches perform much better than the straight trajectory which is expected. However, performing a full survey versus half survey does not provide additional benefit. The full survey does have some trajectories that perform better but in other trajectories it actually performs worse. Therefore, we believe that performing a half survey provides enough positional variance to constrain the localization.

### v3

@@ 3.5.3 Filtering System
Algorithm 2 provides the full outline for the ShadowNav particle filtering algorithm. The algorithm takes as input a Gaussian belief distribution $(\mu_{0},\Sigma_{0})$ assumed for the initial robot position, the number of particles $N_{s}$ to use in the particle filter, and a threshold for the effective number of beliefs $N_{\textrm{eff,thresh}}$ used to trigger resampling (Line 2). The filter is initialized with $N_{s}$ particles drawn from the initial belief distribution and each assigned equal weight (Lines 1-2). Given a new set of crater observations (Line 5), a set of Q-Score measurements is initialized for computing for each individual particle (Line 6). After applying the motion model update to each particle (Line 8), the Q-Score for each updated particle is computed using the procedure from Alg. 1 by comparing against the current measurements (Line 9). As is standard in many particle filtering implementations [43], we note that the particle weights are updated in the in $\log$-domain (Line 13) with a normalization step to ensure non-negative weights (Line 11). Next, the number of effective samples $N_{\textrm{eff}}$ at the current iteration is calculated (Line 15). If $N_{\textrm{eff}}$ is below the threshold $N_{\textrm{eff,thresh}}$, then this is seen as an indication of particle filter “degeneracy”, wherein the weights $\{w_{i}^{t}\}$ collapse around a handful of particles. In such a case, a new set of particles are resampled using the systematic resampling scheme (Line 17).
@@ 4 DATA COLLECTION
In this section, we present details on the datasets generated to validate the efficacy of the proposed ShadowNav approach. We first review the simulation environment used to generated photorealistic stereo data and then review the configuration used to collect nighttime field testing data from Cinder Lakes, Arizona.
@@ 4.1 Simulation
We developed a simulation environment using the Blender software, which has become a popular software framework for generating photorealistic images in a Lunar environment[44, 13]. Within this simulation, the Hapke lighting model [45, 46, 47] was implemented to approximate the Lunar surface reflectance. This model will simulate the “opposition effect” which leads to a focused point of extreme saturation at a location within an image where the camera ray and light source are at zero phase angle. We implemented the Hapke lighting model using “old highland” parameters of the Moon provided in [48]. For our use case, the coherent backscattering opposition effect (CBOE) was left out of our implementation and only the shadow hiding opposition effect (SHOE) was implemented as it dominates most or all lighting calculations.
To obtain a realistic 3D model of the surface geometry, DEMs produced from LROC are utilized.  LROC resolution is typically between $2\text{\,}\mathrm{m}$-$5\text{\,}\mathrm{m}$ which allows for resolving craters of approximately $10\text{\,}\mathrm{m}$. However, the LROC resolution is insufficient for generating smooth surface imagery within simulation. Therefore, the DEMs from LROC are scaled down to be $0.25\text{\,}\mathrm{m}$ resolution before being loaded into Blender. Crater measurements for simulated data in future discussions are based on this scaled resolution. This scaled DEM is imported into Blender and a surface texture comprised of two scales of fractal Brownian motion is added. This texture is a natural noise and simulates surface texture features for stereo algorithms to utilize.
@@ 4.1.1 Simulated Dataset
For simulation, we developed two datasets - one with the aim of testing the efficacy of our crater matching algorithm and a second dataset for testing the overall global localization pipeline. The crater matching dataset was created by selecting four craters of diameters between $5\text{\,}\mathrm{m}$ and $20\text{\,}\mathrm{m}$ from the orbital data. The simulator was run to generate stereo pairs every $1\text{\,}\mathrm{m}$ at four angles around the crater (0, 90, 180, and 270) and at ranges between $5\text{\,}\mathrm{m}$-$25\text{\,}\mathrm{m}$. Additional datasets were created to test sensitivity to the parameters of the hardware configuration:
@@
  • Light offsets of $0.2\text{\,}\mathrm{m}$ below the camera, inline with the camera and $0.2\text{\,}\mathrm{m}$ above the camera
  • Light power ranging from low, medium and high  [^1] As Blender does not accurately model exposure times, the set of “light power” Blender parameters is used to change the brightness of rendered images.
@@ 4.2 Cinder Lakes Apollo Training Area
For the data collection, we performed two types of collections. The first type was geared toward perception tests. This involved moving the data collection rig from approximately $20\text{\,}\mathrm{m}$ away to $5\text{\,}\mathrm{m}$ from the leading crater edge with stops for long exposures roughly every $1\text{\,}\mathrm{m}$. The second type was geared toward localization tests. This entailed moving the data collection rig through the crater field with stops for long exposure approximately every $5\text{\,}\mathrm{m}$-$10\text{\,}\mathrm{m}$. During these stops the rig was directed toward large craters if any were present. In total, in this work, 12 craters ranging from $5\text{\,}\mathrm{m}$-$25\text{\,}\mathrm{m}$ in diameter and three trajectories were collected. The three different trajectories are displayed in Figure 5 which shows the GPS path within Google Earth satellite imagery. Of these, two trajectories were captured in the south site and one in the north site. In order to evaluate longer distances between craters, some experiments added drift error as if more distance had been traveled by artificially adding trajectory locations outside of the crater field at the start of a trajectory.
@@ 5.1 Stereo
Accurate stereo is required to be able to effectively detect and match crater rims within our approach. For detection, outliers in range values or stereo holes could lead to crater rim false positives. Furthermore, inaccurate stereo ranges could lead to poor projections into the orbital map which could cause correctly identified crater rims in image-space to match poorly to orbital landmark craters. To evaluate stereo performance, we created a simulated dataset of craters with ground truth ranges which was detailed in Section 4.1.
To evaluate stereo performance, three metrics based on [49] binned across ground truth range values are used. These metrics are:
@@
We note that a crucial difference between the metrics used in [49] is that range is used rather than disparity as range is what ultimately has the impact on crater matching.
To validate the impact of hardware system parameters such as light location, camera height, and light power, these parameters were swept and stereo performance evaluated. Light power is used as a proxy for exposure time and light brightness as these are not modelled well in Blender. High light power corresponds to the near range of the image being fully exposed (pixel value of 255). A zero-mean 1.2 pixel intensity sigma Gaussian noise was added to all images to provide additional realism for nighttime imagery. The results of this analysis are in Table 1. In these results, we observe that the optimal configuration for stereo is with a $2.5\text{\,}\mathrm{m}$ camera off ground, light offset $0.2\text{\,}\mathrm{m}$ below the camera, and high light power. Medium light power at a camera height of $2.5\text{\,}\mathrm{m}$, wherein the near range is bright but few pixels are fully exposed, demonstrates improved performance within $30\text{\,}\mathrm{m}$ but at the slight cost of performance beyond $30\text{\,}\mathrm{m}$. Moreover, although this configuration has robustness to camera height ranges, we observe that a camera height of $1.5\text{\,}\mathrm{m}$ performs slightly worse than at $2.5\text{\,}\mathrm{m}$ and this is attributed to cameras closer to the ground losing vertical resolution with range. Finally, a key takeaway is that placing the camera inline with the camera significantly degrades performance and placing the light above the camera renders extremely poor performance . We found that placing the light below the camera minimizes the worst impacts of the opposition effect and provides some shadowing on the surface for more visible features.
@@ 5.2 Crater Detection and Matching
To evaluate crater detection and matching, we used the real data collected at Cinder Lakes detailed in Section 4.2. Additional simulated data was collected to evaluate stereo and positional errors as detailed in Section 4.1. We use the percent of the ground truth front rim detected as a metric to evaluate detection performance:
@@
,where $px_{\textrm{gt}_{\textrm{matched}}}$ is the number of ground truth crater points in world frame that have a detection that is closest to that ground truth point and $px_{\textrm{gt}}$ is the total number of all ground truth points. This detection does not consider false positives. To evaluate crater matching performance, we use the Q-Score metric detailed in Section 3.5.2 to evaluate how closely a perception sample aligns to orbital ground truth considering both true detections and false positives.
SGBM stereo and discontinuity detection were run across the Cinder Lakes crater dataset to evaluate crater detection and matching performance. Figure 6 shows the results with respect to percent of front rim detected and Q-Scores versus range and crater diameter. From these results, we observe that nearly all craters have some level of detection between $5\text{\,}\mathrm{m}$-$15\text{\,}\mathrm{m}$. However, at $15\text{\,}\mathrm{m}$ the $7\text{\,}\mathrm{m}$-crater is no longer detected. Furthermore, the $25\text{\,}\mathrm{m}$ has worse results across most of the set of camera distances from the crater rim. The reduction in performance for smaller and larger craters is expected with our current approach. A larger crater diameter requires stereo to obtain matches at farther ranges to observe the leading and trailing edge of a discontinuity. A smaller crater diameter limits the size of a discontinuity, which in turns limits the ability to detect it with few false positives. Overall, we observe robust filter performance, with most samples containing a Q-Score of 0.4 or better which means that the average pixel detection is within $2\text{\,}\mathrm{m}$ of the crater rim and indicates low false positives and high accuracy.
@@
Figure 7: Crater detection and matching results with an without Gaussian noise added to simulated images on data with a camera height of 2.5m a light offset of 20cm below camera and high exposure. For images with noise added, 0 mean and 1.2 pixel intensity sigma were used. (a) Q-Score no noise (b) Percent detected no noise (c) Q-Score 1.2 sigma noise (d) Percent detected 1.2 sigma noise.
To further evaluate crater detection, simulated data is evaluated first by comparing crater rim detection and matching with and without Gaussian noise added. Figure 7 contains the results of comparing noise free versus images with zero mean, 1.2 sigma pixel intensity Gaussian noise added. From these results, it is observed that the addition of noise has minimal impact on the performance of crater detection. Compared with the Cinder Lakes results, however, detection performance was a bit worse at both close to $5\text{\,}\mathrm{m}$ to the crater and beyond $15\text{\,}\mathrm{m}$ from crater rim though overall similar. The larger difference in performance is observed in the Q-Score results where there is a more significant decrease in performance beyond $15\text{\,}\mathrm{m}$. However, for both real and simulated craters with diameters between $7.5\text{\,}\mathrm{-}$$20\text{\,}\mathrm{m}$ Q-Score results are near or above 0.4 within $15\text{\,}\mathrm{m}$. Potential sources of the performance difference between real and simulated data include the simulation containing images with a wider field-of-view, a $2.5\text{\,}\mathrm{m}$ camera height compared to $1.5\text{\,}\mathrm{m}$ camera height for Cinder Lakes data, and the inherent simulation versus real differences in lighting, camera exposure handling, and surface textures. Using the higher camera height within simulation could reduce the discontinuity distance and explain the performance difference within $7\text{\,}\mathrm{m}$ to the front of the crater rim. The larger field-of-view camera within simulation would allow for more potential false positives to be captured as evidenced by strong crater rim percent detection but degraded Q-Scores at $15\text{\,}\mathrm{m}$-$20\text{\,}\mathrm{m}$ ranges to crater rims.
@@
Figure 9: Crater detection and matching results with positional error on simulated data with a camera height of 2.5m a light offset of 20cm below camera, high brightness, and 1.2 pixel intensity sigma Gaussian noise to image (a) Q-Score 5m error (b) Percent detected 5m error (c) Q-Score 10m error (d) Percent detected 10m error.
One of the limitations of the above analysis is that all of the craters were mostly centered within the image frame. This centering will only be possible with both an accurate orientation and position knowledge, However, upon approaching a crater, there is an expected error in absolute position due to the drift of relative localization. To evaluate how these positional errors affect detection performance, data was rendered with a positional offset from the position from which the ideal heading was computed. This data is described in more detail in Section 4.1. Using this data, Figure 9 shows the results of computing the Q-Scores and percents of crater rims detected with a $5\text{\,}\mathrm{m}$ and $10\text{\,}\mathrm{m}$ horizontal shift. Only horizontal errors are plotted because forward/backward errors are captured as a shift along the x-axis in Figure 7. From Figure 9, we observe that a $5\text{\,}\mathrm{m}$ horizontal offset contains some reduction in crater matching performance but percent of crater rim detected maintains similar performance without a position offset. However, the performance of crater matching and detection degrades at less than $10\text{\,}\mathrm{m}$ range due to the crater not being captured within the field-of-view of the camera. Crater matching performance still degrades around $15\text{\,}\mathrm{m}$, but the degradation is similar to the results without a positional offset. Overall crater detection and matching demonstrates strong performance within the $5\text{\,}\mathrm{m}$-$15\text{\,}\mathrm{m}$ range with potential for performance out to $20\text{\,}\mathrm{m}$ detection range.
@@
5.3 Absolute Localization
Instantaneous perception measurements demonstrated strong performance in the previous section, but also presented some issues. Through the use of a particle filter approach outlined in Section 3.5, multiple perception measurements can be fused overtime to correct absolute position after it has drifted. Both simulation and Cinder Lakes data is used to evaluate the performance of the absolute localization.
5.3.1 Simulation